### v1

# MobileCLIP: Fast Image-Text Models through Multi-Modal Reinforced Training

Abstract
Contrastive pretraining of image-text foundation models, such as CLIP, demonstrated excellent zero-shot performance and improved robustness on a wide range of downstream tasks. However, these models utilize large transformer-based encoders with significant memory and latency overhead which pose challenges for deployment on mobile devices. In this work, we introduce MobileCLIP – a new family of efficient image-text models optimized for runtime performance along with a novel and efficient training approach, namely multi-modal reinforced training. The proposed training approach leverages knowledge transfer from an image captioning model and an ensemble of strong CLIP encoders to improve the accuracy of efficient models. Our approach avoids train-time compute overhead by storing the additional knowledge in a reinforced dataset. MobileCLIP sets a new state-of-the-art latency-accuracy tradeoff for zero-shot classification and retrieval tasks on several datasets. Our MobileCLIP-S2 variant is 2.3$\times$ faster while more accurate compared to previous best CLIP model based on ViT-B/16. We further demonstrate the effectiveness of our multi-modal reinforced training by training a CLIP model based on ViT-B/16 image backbone and achieving +2.9% average performance improvement on 38 evaluation benchmarks compared to the previous best. Moreover, we show that the proposed approach achieves 10$\times$-1000$\times$ improved learning efficiency when compared with non-reinforced CLIP training.
Figure 1: MobileCLIP models are fast and accurate. Comparison of publicly available CLIP models with MobileCLIP trained on our DataCompDR dataset. Latency is measured on iPhone12 Pro Max.
Figure 2: DataCompDR dataset improves all metrics. Zero-shot performance of CLIP models with ViT-B/16 image encoder.
1 Introduction
Large image-text foundation models, such as CLIP [46], have demonstrated excellent zero-shot performance and improved robustness [15] across a wide range of downstream tasks [29]. However, deploying these models on mobile devices is challenging due to their large size and high latency.
Our goal is to design a new family of aligned image-text encoders suitable for mobile devices. There are two main challenges towards realizing this goal. First, there is a tradeoff between runtime performance (e.g., latency) and the accuracy of different architectures, therefore we should be able to quickly and thoroughly analyze different architectural designs. Large-scale training of CLIP models is computationally expensive, hindering rapid development and exploration of efficient architecture design. On the other hand, standard multi-modal contrastive learning [46] at small-scale results in poor accuracies, which do not provide a useful signal to guide architecture design choices. Second, reduced capacity of smaller architectures leads to subpar accuracy that can be improved with a better training method.
To overcome these challenges, we develop a novel training approach based on the dataset reinforcement method [14]: i) reinforce a dataset once with additional information, and ii) use the reinforced dataset several times for experimentation. For a given compute budget, training with the reinforced dataset results in improved accuracy compared to the original dataset. We propose a multi-modal variant of dataset reinforcement for training efficient CLIP models. Specifically, we reinforce the image-text DataComp [18] dataset by adding synthetic captions and embeddings from a strong ensemble of pretrained CLIP models (Fig. 3), obtaining DataCompDR. We introduce two variants of our reinforced dataset, DataCompDR-12M suited for rapid iteration on efficient model design and DataCompDR-1B for best large-scale training performance.
Training with DataCompDR shows significant learning efficiency improvement compared to the standard CLIP training. For example, with a single node of 8$\times$×A100 GPUs, we achieve 61.7% zero-shot classification on ImageNet-val [8] in approximately one day when training a ViT-B/16 [12] based CLIP from scratch on DataCompDR-12M. Training with DataCompDR-1B sets new state-of-the-art performance on several metrics (Fig. 2) while still using a fraction of the training compute budget compared to previous works.
Utilizing DataCompDR, we explored the design space and obtained a new family of mobile-friendly aligned image-text encoders called MobileCLIP with a better latency-accuracy tradeoff compared to the previous works (Fig. 1). We exploit several architectural design techniques to obtain efficient image and text encoders, including structural reparametrization [9, 20, 10, 11, 60] and convolutional token mixing [61]. MobileCLIP includes S0, S1, S2, and B variants covering various sizes and latencies for different mobile applications. Our fastest variant, MobileCLIP-S0, is approximately 5$\times$ faster and 3$\times$ smaller than the standard OpenAI ViT-B/16 CLIP model [46], but has the same average accuracy. Our contributions are as follows:
  • We design a new family of mobile-friendly CLIP models, MobileCLIP. Variants of MobileCLIP use hybrid CNN-transformer architectures with structural reparametrization in image and text encoders to reduce the size and latency.
  • We introduce multi-modal reinforced training, a novel training strategy that incorporates knowledge transfer from a pre-trained image captioning model and an ensemble of strong CLIP models to improve learning efficiency.
  • We introduce two variants of our reinforced datasets: DataCompDR-12M and DataCompDR-1B. Using DataCompDR, we demonstrate 10x-1000x learning efficiency in comparison to DataComp.
  • MobileCLIP family obtains state-of-the-art latency-accuracy tradeoff on zero-shot tasks, including marking a new best ViT-B/16 based CLIP model.
2 Related Work
Efficient learning for CLIP.
One can improve learning efficiency through utilizing an enhanced training objective. Examples include image masking [70, 17, 54, 36], unimodal self-supervision [42, 34], fine-grained image-text alignment [71], contrastive learning in image-text-label space [68], and pairwise Sigmoid loss [76]. Recently, CLIPA [33] proposed training at multi-resolutions for cost-effective CLIP training. These methods are complementary to our proposed method and can be exploited for further improvements.
CLIP training dataset is often comprising noisy image-text pairs obtained at web-scale. Since the original CLIP model [46], several works have demonstrated improved results on large-scale and filtered datasets [50, 51, 18, 76, 16]. Complementary to data collection and filtering, recent works show that using visually enriched synthetic captions generated from a pretrained captioning model, along with real captions, can improve the quality of CLIP models [69, 44, 31]. Our proposed reinforced multi-modal dataset also benefits from synthetically generated captions, which we show are crucial for improved learning efficiency.
Previous works explored extending unimodal knowledge distillation [25] to vision-language models. DIME-FM [55] proposes using in-domain unimodal data for distillation with a focus on zero-shot classification. TinyCLIP [67] trains compact CLIP models via cross-modal affinity mimicking and weight inheritance. Multi-modal distillation is also explored in setups where the student is a fused vision-language model for specific tasks [64, 30, 63]. Our proposed multi-modal reinforced training also includes cross-modal affinity mimicking [67] toward targets that are added to our reinforced datasets. Further, we extend unimodal model ensembling [32, 45] to multimodal setup, and store targets obtained from an ensemble of CLIP models.
Offline knowledge distillation methods [53, 75, 14] have been proposed recently to mitigate the training-time overhead cost due to running large teacher models. We extend the dataset reinforcement strategy [14] to the multi-modal setup of CLIP. Our proposed reinforced multi-modal datasets result in significant accuracy improvement without adding a training-time computational overhead.
Efficient architectures for CLIP.
Recently there have been a wide range of architectures that have shown great promise for accomplishing vision tasks on resource constraint devices. These architectures can be broadly classified into purely convolutional [22, 40, 47, 11, 27, 49, 26, 60], transformer based [12, 58, 39] and convolution-transformer hybrids like [35, 37, 61, 21, 52, 43]. Similarly there are transformer based [62] and convolution-transformer hybrids like [66, 19] for text encoding. There have been works like [67], that prune ViT architectures to obtain smaller and faster CLIP models or works like [3] that reduce image-text tokens for faster inference of vision-language models. These models can still be quite large and inefficient to be deployed on a mobile device. In our work, we introduce an improved convolution-transformer hybrid architecture for both vision and text modalities, that improve over recent state-of-the-art like [37, 52, 43, 21]. The optimizations introduced in  [67, 3] can be used to further improve efficiency of our models.
3 Multi-Modal Reinforced Training
Our multi-modal reinforced training leverages knowledge transfer from an image captioning model and a strong ensemble of pretrained CLIP models for training the target model. It consists of two main components: i) leveraging the knowledge of an image captioning model via synthetic captions, and ii) knowledge distillation of image-text alignments from an ensemble of strong pre-trained CLIP models. We follow the dataset reinforcement strategy of [14] and store the additional knowledge (synthetic captions and teacher embeddings) in the dataset (see Fig. 3), thereby avoiding any additional training time computational overhead such as evaluating the captioning model or the ensemble teacher. The proposed training strategy results in significant improvement in learning efficiency, i.e., reaching to certain target performance with less training budget and fewer samples.
[Uncaptioned image] Figure 3: Illustration of multi-modal dataset reinforcement with one image augmentation and one synthetic caption. In practice, we use multiple image augmentations and synthetic captions.
[Uncaptioned image] Figure 4: Architecture of convolutional and reparameterizable blocks, called Text-RepMixer used in MobileCLIP’s text encoder - MCt.
3.1 Dataset Reinforcement
Synthetic captions.
Image-text datasets used to train CLIP models are mostly sourced from the web, which is inherently noisy. Recent efforts such as DataComp [18] and data filtering networks [16] improve the quality of web-sourced datasets by using extensive filtering mechanisms. While these filtered datasets have lower noise, the captions may still not be descriptive enough. In order to boost the visual descriptiveness of the captions we use the popular CoCa [73] model and generate multiple synthetic captions ${{\bm{x}}}_{\text{syn}}^{(i,s)}$ for each image ${{\bm{x}}}_{\text{img}}^{(i)}$ (see Fig. 3a). Ablations on the number of synthetic captions generated per image are provided in Sec. 5.1. Figure 5 shows some examples of synthetic captions generated by the CoCa model. Real captions in comparison to synthetic captions are generally more specific but noisier. We show (Tab. 1(a)) a combination of both real and synthetic captions is crucial to obtain best zero-shot retrieval and classification performance.
Image augmentations.
For each image ${{\bm{x}}}_{\text{img}}^{(i)}$, we generate multiple augmented images ${\hat{{\bm{x}}}}_{\text{img}}^{(i,j)}$ using a parametrized augmentation function $\mathcal{A}$:
${\hat{{\bm{x}}}}_{\text{img}}^{(i,j)}=\mathcal{A}({{\bm{x}}}_{\text{img}}^{(i)};{{\bm{a}}}^{(i,j)})\,,$ (1)
where ${{\bm{a}}}^{(i,j)}$ are the augmentation parameters that are sufficient to reproduce ${\hat{{\bm{x}}}}_{\text{img}}^{(i,j)}$ from ${{\bm{x}}}_{\text{img}}^{(i)}$ (see Fig. 3a). Ablations on the number and different kinds of augmentations used per image are provided in Tabs. 2(a) and 11, respectively.
Ensemble teacher.
Model ensembling is a widely used technique for creating a stronger model from a set of independently trained ones [32, 45]. We extend this technique to multi-modal setup and use an ensemble of $K$ CLIP models as a strong teacher (see Sec. 5.1 for our teacher ablations). We compute the feature embeddings of these models for augmented images ${\hat{{\bm{x}}}}_{\text{img}}^{(i,j)}$ and synthetic captions ${{\bm{x}}}_{\text{syn}}^{(i,s)}$ obtaining $d_{k}$-dimensional vectors ${\psi_{\text{img}}^{(i,j,k)}}$ and ${\psi_{\text{syn}}^{(i,s,k)}}$ for the $k$-th teacher model. We also compute the teacher embeddings ${\psi_{\text{txt}}^{(i,k)}}$ of the ground-truth captions ${{\bm{x}}}_{\text{txt}}^{(i)}$ (see Fig. 3b).
Reinforced dataset.
We store the image augmentation parameters ${{\bm{a}}}^{(i,j)}$, synthetic captions ${{\bm{x}}}_{\text{syn}}^{(i,s)}$, feature embeddings ${\psi_{\text{img}}^{(i,j,k)}}$, ${\psi_{\text{syn}}^{(i,s,k)}}$ and ${\psi_{\text{txt}}^{(i,k)}}$ of the CLIP teachers as additional knowledge in the dataset along with the original image ${{\bm{x}}}_{\text{img}}^{(i)}$ and caption ${{\bm{x}}}_{\text{txt}}^{(i)}$ (see Fig. 3c). Note that dataset reinforcement is a one-time cost that is amortized by several efficient model training and experimentation.
3.2 Training
Loss function.
Intuitively, our loss function distills the affinity matrix between image-text pairs from multiple image-text teacher encoders into student image-text encoders. Let $\mathcal{B}$ denote a batch of $b$ (image, text) pairs and ${\Psi_{\text{img}}^{(k)}},{\Psi_{\text{txt}}^{(k)}}\in\mathcal{R}^{b\times d_{k}}$ the matrices of $d_{k}$-dimensional image and text embeddings, respectively, of the $k$-th model in the teacher ensemble for batch $\mathcal{B}$. Correspondingly, we denote the image and text embedding matrices of the target model by ${\varPhi_{\text{img}}},{\varPhi_{\text{txt}}}\in\mathcal{R}^{b\times d}$. For given ${\bm{U}}$ and ${\bm{V}}$ matrices, let ${\mathcal{S}}_{\tau}({\bm{U}},{\bm{V}})\in\mathcal{R}^{b\times b}$ denote their similarity matrix obtained by applying row-wise Softmax operation to ${\bm{U}}{\bm{V}}^{\top}/\tau$, where $\tau$ is a temperature parameter. Our training loss consists of two components, the standard CLIP [46] loss ${\mathcal{L}_{\text{CLIP}}}(\mathcal{B})$ and a knowledge distillation loss ${\mathcal{L}_{\text{Distill}}}(\mathcal{B})$:
${\mathcal{L}_{\text{Total}}}(\mathcal{B})$ $=\lambda{\mathcal{L}_{\text{CLIP}}}(\mathcal{B})+(1-\lambda){\mathcal{L}_{\text{Distill}}}(\mathcal{B}),$ (2)
${\mathcal{L}_{\text{Distill}}}(\mathcal{B})$ $=\frac{1}{2}{\mathcal{L}_{\text{Distill}}^{\text{I2T}}}(\mathcal{B})+\frac{1}{2}{\mathcal{L}_{\text{Distill}}^{\text{T2I}}}(\mathcal{B}),$
${\mathcal{L}_{\text{Distill}}^{\text{I2T}}}(\mathcal{B})$ $=\frac{1}{bK}\sum_{k=1}^{K}\text{KL}({\mathcal{S}}_{\tau_{k}}({\Psi_{\text{img}}^{(k)}},{\Psi_{\text{txt}}^{(k)}})\|{\mathcal{S}}_{\widehat{\tau}}({\varPhi_{\text{img}}},{\varPhi_{\text{txt}}})),$
where KL denotes Kullback-Leibler divergence, ${\mathcal{L}_{\text{Distill}}^{\text{T2I}}}$ is computed by swapping the text and image embedding terms of ${\mathcal{L}_{\text{Distill}}^{\text{I2T}}}$, and $\lambda$ is a tradeoff parameter.
Figure 5: Real vs synthetic captions.
Efficient training.
Training on the reinforced dataset is as simple as modifying the data loader and loss function to exploit additional knowledge stored in the dataset and has the same training cost as standard CLIP training (see Tab. 2(d)). For every sample, we read the image ${{\bm{x}}}_{\text{img}}^{(i)}$ and the corresponding ground-truth caption ${{\bm{x}}}_{\text{txt}}^{(i)}$ from the dataset. Then, we randomly load one of stored augmentation parameters ${{\bm{a}}}^{(i,j)}$ and reproduce the augmented image ${\hat{{\bm{x}}}}_{\text{img}}^{(i,j)}$. We also randomly load one of synthetic captions ${{\bm{x}}}_{\text{syn}}^{(i,s)}$. Finally, we read the stored embeddings, ${\psi_{\text{img}}^{(i,j,k)}}$, ${\psi_{\text{syn}}^{(i,s,k)}}$, and ${\psi_{\text{txt}}^{(i,k)}}$, corresponding to the $K$ teacher models.
Using this loaded data, we construct two data batches, $\mathcal{B}_{\text{real}}$ corresponding to (augmented image, real caption) pairs and $\mathcal{B}_{\text{syn}}$ corresponding to (augmented image, synthetic caption) pairs, and compute our training loss in Eq. 2 separately on $\mathcal{B}_{\text{real}}$ and $\mathcal{B}_{\text{syn}}$. Our final loss is given by
$\sum_{\mathcal{B}\in\{B_{\text{real}},B_{\text{syn}}\}}{\mathcal{L}_{\text{Total}}}(\mathcal{B}).$ (3)
Note that we can compute the total loss after a forward pass of the student model without any extra teacher-related computations since the teacher embeddings required to compute the distillation loss are readily available as part of the dataset.
4 Architecture
4.1 Text Encoder
CLIP [46] model paired the vision transformer with a classical transformer comprising of self-attention layers for text encoding. While this model is effective, smaller and more efficient models are preferred for mobile deployment. Recently, works like [66] have shown that convolutions can be just as effective for text encoding. In contrast, we found that purely convolutional architectures significantly underperform their transformer counterparts. Instead of using a fully convolutional architecture for text encoding, we introduce a hybrid text encoder which makes use of 1-D convolutions and self-attention layers. We design variants of MobileCLIP using both transformer and hybrid text encoders.
For hybrid text encoder, we introduce Text-RepMixer, a convolutional token mixer that decouples train-time and inference-time architectures. Text-RepMixer is inspired by reparameterizable convolutional token mixing (RepMixer) introduced in [61]. At inference, skip connections are reparameterized. The architecture is shown in Fig. 4. For Feed-Forward Network (FFN) blocks, we augment linear layers with an additional depthwise 1-D convolution of similar kernel dimensions as the token mixer, to obtain ConvFFN blocks. This structure is similar to the convolutional blocks used in [19], the main difference being the use of batchnorm and the ability to fold it with the succeeding depthwise 1-D convolutional layer for efficient inference. For all depthwise 1-D convolutions, we use a kernel size of 11. This was chosen based on accuracy-latency tradeoff, full ablation can be found in Appendix E. In order to find the optimal design for our hybrid text encoder, we start with a purely convolutional text encoder and start replacing convolutional blocks systematically with self-attention layers (see Tab. 3(b)).
4.2 Image Encoder
Recent works have shown the efficacy of hybrid vision transformer for learning good visual representations. For MobileCLIP, we introduce an improved hybrid vision transformer called MCi based on the recent FastViT [61] architecture with certain key differences explained below.
In FastViT, an MLP expansion ratio of 4.0 is used for FFN blocks. Recent works like [38, 67] exposed the significant amount of redundancy in linear layers of FFN block. To improve parameter efficiency, we simply lower the expansion ratio to 3.0 and increase the depth of the architecture. By doing so, we retain the same number of parameters in the image encoder. The stage configuration for the three variants are described in Tab. 3(a). MCi0 has similar stage configuration as [60]. MCi1, is a deeper version of MCi0 and MCi2 is a wider version of MCi1. The stage compute ratios in our variants are similar to [60]. We find that this design has a minimal impact on latency, but a good improvement in capacity of the model as reflected in the downstream task performance. When trained from scratch on ImageNet dataset for image classification task, MCi2 attains the same top-1 accuracy of 84.5% as FastViT [61] (previous state-of-the-art hybrid vision transformer), while being 15% smaller and 14.3% faster. More details can be found in Appendix A.
5 Experiments
In this section, we present our experimental setup, ablations on our proposed method and fast MobileCLIP architectures, and results.
Evaluation.
We evaluate image-text models using the evaluation benchmark of DataComp [18]. Specifically, we report zero-shot classification on the ImageNet validation set [8], and its distribution shifts including ImageNet-V2 [48], ImageNet-A [24], ImageNet-O [24], ImageNet-R [23], and ObjectNet [1], which we report their average as IN-Shift. For zero-shot image-text retrieval, we report recall@1 on MSCOCO [5] and Flickr30k [72] datasets. Further, we report average performance on all 38 datasets in DataComp evaluations. We also evaluate our models on Visual Genome Relation, Visual Genome Attributes, Flickr30k-Order and COCO-Order datasets which are part of the recent Attribute, Relation and Order (ARO) benchmark [74]. In the remainder, IN-val refers to zero-shot accuracy on ImageNet validation set and Flickr30k refers to average zero-shot recall@1 for image-text and text-image retrieval. All reported metrics are obtained without any fine-tuning.
Training setup.
We have two setups for ablations and large-scale experiments. For ablations, we train on datasets with 12.8M image-text pairs using a global batch size of 8,192 and 8$\times$×NVIDIA-A100-80GB GPUs for 30-45k iterations. For large-scale training, we use a global batch size of 65,536 with 256$\times$×A100 GPUs for 200k iterations. All models are trained from scratch (see details in Appendix A).
Dataset.
We train on the image-text dataset of DataComp dataset [18]. We use the Bestpool filtered subset of 1.28B samples that provides them with best performance at the largest dataset scale. We refer to this set as DataComp-1B. For fast experimentation, we create a fixed subset of 12.8M uniformly sampled pairs which we call DataComp-12M. DataComp-12M was not studied in [18] but in our experiments, we observed that DataComp-12M consistently achieves better performance compared with the Bestpool subset of DataComp-medium with comparable samples.
DataCompDR: Reinforced DataComp.
We reinforce the DataComp dataset using our multi-modal dataset reinforcement strategy. In particular, we create DataCompDR-1B and DataCompDR-12M by reinforcing DataComp-1B and DataCompDR-12M. We have a one-time generation process, the cost of which is amortized over multiple architectures and extensive ablations. We generate 5 synthetic captions per image using the coca_ViT-L-14 model in OpenCLIP [28], and strong random image augmentations (10 for DataCompDR-1B and 30 for DataCompDR-12M). We compute embeddings of an ensemble of two strong teachers (ViT-L-14 with pretrained weights datacomp_xl_s13b_b90k and openai in OpenCLIP) on augmented images as well as real and synthetic captions. Embeddings are 1536-D concatenations of 2$\times$768-D vectors. We store all reinforcements using lossless compression and BFloat16. We analyze all of our choices in Sec. 5.1.
MobileCLIP architectures.
Our MobileCLIP architectures are formed as pairs of MCi:MCt architectures. In particular, we create 3 small variants MobileCLIP-S0 (MCi0:MCt), MobileCLIP-S1 (MCi1:Base), and MobileCLIP-S2 (MCi2:Base), where Base is a 12-layer Transformer similar to the text-encoder of ViT-B/16 based CLIP [46]. We also train a standard pair of ViT-B/16:Base and refer to our trained model as MobileCLIP-B.
Benchmarking latency.
To measure latency, we use the input sizes corresponding to the respective methods. For iPhone latency measurements, we export the models using Core ML Tools (v7.0) [57] and run it on iPhone12 Pro Max with iOS 17.0.3. Batch size is set to 1 for all the models. We follow the same protocol as described in [60].
5.1 Ablation Studies
$\lambda$
Syn. Captions
Strong Aug.
Ens. Teacher
IN-val
Flickr30k
0
44.5
41.8
0
51.9
69.3
1
54.5
66.1
1
59.3
70.5
1
61.7
72.0
0.7
60.7
74.2
Table 1: Summary of ablations. We train on DataCompDR-12M for 30k iterations. All ablations are on ViT-B/16:Base. We highlight our main choices with blue and alternative tradeoffs with gray. We underline numbers within $0.5\%$ of the maximum.
In this section, we analyze the effect of each component in our training and architecture. Unless otherwise stated, we use ViT-B/16:Base encoders trained on DataComp-12M for 30k iterations with global batch size of 8k ($\sim$20 epochs). Table 1 summarizes the analysis of our training.
Strong image augmentations.
In contrast to uni-modal supervised and self-supervised methods for vision where strong augmentations are used [13, 59], CLIP training recipes [46] often use light image augmentations to avoid image-text mis-alignment. However, several works [45, 14, 2] demonstrated the efficacy of strong augmentations in a distillation setup. In Tab. 1 we show that strong image augmentations improve distillation performance (+4.8% on IN-val and +4.4% on Flickr30k). We provide detailed ablation on the effect of image augmentations in Appendix B.
Synthetic captions.
Similar to image augmentations, synthetic captions (or caption augmentations) can further improve the performance of CLIP models, particularly on image-text retrieval. For regular CLIP training ($\lambda=0$), we observe in Tab. 1 that including batches with both synthetic and real captions results in +7.4% on IN-val and +27.5% on Flickr30k performance improvements. In Tab. 1(a), we observe a similar trend for CLIP training with distillation loss only ($\lambda=1$). In Tab. 1(b), we analyze the effect of $\lambda$ and observe a tradeoff where $\lambda=1.0$ is optimal for IN-val while $\lambda=0.7$ is optimal for Flickr30k. Prior work that exploit synthetic captions primarily focus on improved retrieval [31, 69] while distillation works focus on zero-shot classification [55]. In our large-scale experiments, we balance the tradeoff for MobileCLIP-B using $\lambda=0.75$ and use $\lambda=1.0$ for our small variants.
Ensemble teacher.
We find that using an ensemble of strong CLIP models as a teacher in our multi-modal reinforced training is crucial to achieving +2.4% IN-val improvement (Tab. 1). We also observe that the most accurate models are not the best teachers. See Appendix C for a comprehensive analysis of different teacher models.
$\mathcal{B}\in$$\{\mathcal{B}_{\text{real}}\}$$\{\mathcal{B}_{\text{syn}}\}$$\{\mathcal{B}_{\text{real}}$ or $\mathcal{B}_{\text{syn}}\}$$\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$
IN-val 56.4 49.8 57.3 61.7
Flickr30k 57.0 72.2 68.6 72.0
(a) Real vs synthetic sampling in Eq. 3 ($\lambda=1.0$).
$\lambda$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
IN-val 54.4 56.3 57.4 58.2 59.5 60.3 60.7 61.5 61.6 61.7
Flickr30k 71.4 71.5 71.8 72.2 73.8 73.6 74.2 73.1 73.2 72.0
(b) Ablation on the loss coefficient ($\lambda$) in Eq. 2.
Table 2: Ablation on the loss. The tradeoff between IN-val and Flickr30k is controlled by the synthetic sampling and loss coefficient. We train for 30k iterations.
Num. Aug. 1 2 5 10 15 20 25 30
IN-val 60.63 63.27 64.81 64.74 64.49 64.92 64.78 64.74
Flickr30k 69.61 71.74 74.76 74.46 73.90 74.29 73.27 75.66
(a) Effect of the number of augmentations.
Num. Caps. 0 1 2 3 4 5
IN-val 60.67 64.88 65.19 65.19 64.81 64.74
Flickr30k 62.26 73.82 74.27 73.91 74.07 75.66
(b) Effect of the number of synthetic captions.
Dataset Image Text Syn.
Aug. Params
Text Emb.
Image Emb.
Size (TBs)
DataComp-12M
0.9
0.9
DataCompDR-12M
5+1
30
1.9
DataComp-1B
90
DataCompDR-1B
5+1
10
140
(c) Total storage for samples stored in individual Pickle Gzip files and BFloat16 embeddings. $+1$ refers to the ground-truth caption. For further size reductions see Tab. 14.
Dataset
$\mathcal{B}\in$
${\mathcal{L}_{\text{CLIP}}}$
${\mathcal{L}_{\text{Distill}}}$
Stored Syn. Caption
Stored Embeddings
Time (hours)
DataComp-12M
$\{\mathcal{B}_{\text{real}}\}$
1.3
$\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$
21.1
$\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$
4.1
DataCompDR-12M
$\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$
1.3
(d) Training time per epoch (12.8M samples) on 8×8$\times$8 ×A100-80GB.
Table 3: Ablations on storage/cost. Training on DataCompDR has no time overhead. We train for 45k iterations ($\sim$30 epochs).
Variant $\{C_{1},C_{2},C_{3},C_{4}\}$$\{L_{1},L_{2},L_{3},L_{4}\}$
MCi0 $\{64,128,256,512\}$$\{2,6,10,2\}$
MCi1 $\{64,128,256,512\}$$\{4,12,20,4\}$
MCi2 $\{80,160,320,640\}$$\{4,12,24,4\}$
(a) Configurations of MCi.
Num. Self-attn. 6 4 2 1 0
Num Params. (M) 44.5 42.4 40.4 39.3 38.3
Latency (ms) 1.9 1.6 1.4 1.3 1.2
IN-val 60.9 60.8 60.2 60.0 57.9
(b) Effect of the number of self-attention layers in MCt.
Table 4: Ablation on architecture. We train for 30k iterations.
Number of image augmentations and synthetic captions.
We generate multiple image augmentations and synthetic captions and store them efficiently along with the teacher embeddings. We investigate the effectiveness of the number of augmentations and synthetic captions in Tabs. 2(a) and 2(b). We train models with up to 30 image augmentations and 5 synthetic captions for 45k iterations ($\sim$30 epochs). We observe that the performance nearly saturates at 5 augmentations and 2 synthetic captions suggesting each augmentation can be reused multiple times before the added knowledge is fully learned by the model. When needed, fewer augmentations and synthetic captions can help reduce the generation time and storage overhead. For maximal performance, we reinforce DataCompDR-12M and DataCompDR-1B with 10 and 30 augmentations, respectively, and 5 synthetic captions.
Training time.
A major advantage of reinforced training is the minimal time difference with non-reinforced training. We provide the wall-clock times in Tab. 2(d) for regular CLIP training as well as training with online distillation and a caption generator. We measure the time for training on one epoch of DataCompDR-12M on a single node with 8$\times$ A100-80GB GPUs. An epoch takes 1562 iterations with global batch size 8192 on DataCompDR-12M. Without any dataset reinforcement, training is 16$\times$ slower while with partial reinforcements of synthetic captions it is 3$\times$ slower.
Storage size.
We report the storage requirements for our reinforced datasets compared with the original DataComp dataset. In general, the storage size of datasets depends on the file format and the tradeoff between load time and the compression rate. We report the storage size of one file per image-text pair. If present, we store all corresponding reinforcements in the same file. We store files in the Pickle format and compress each file with Gzip compression. The image-text embeddings are saved in BFloat16. We report the total storage size for 12.8M samples of DataCompDR-12M and 1.28B samples of DataCompDR-1B in Tab. 2(c). We provide analysis on additional size reductions in Appendix D and verify that using BFloat16 does not impact the accuracy. For minimal storage overhead, we recommend 5 augmentations/synthetic captions for 30 epochs on DataCompDR-12M and 2 for 10 epochs on DataCompDR-1B which are based on our ablations in Tabs. 2(a) and 2(b).
Hybrid text encoder.
We ablate over the number of Text-RepMixer blocks that can effectively replace self-attention layers with negligible impact on zero-shot performance. For this ablation, we choose a 6-layer purely convolutional text encoder and systematically introduce self-attention layers in the middle. From Tab. 3(b), we find that even introducing a single self-attention layer substantially improves the zero-shot performance. The best tradeoff is with 2 blocks of Text-RepMixer and 4 blocks of self-attention layers. This variant, MCt, obtains similar performance as the pure transformer variant, while being 5% smaller and 15.8% faster.
5.2 Small Scale Regime
In Tab. 5, we compare methods trained on datasets with 12-20M samples, a relatively small range for fast exploration (e.g., architecture search). MobileCLIP-B trained on DataCompDR-12M with less than 370M samples significantly outperforms all other methods with up to 4$\times$ longer training. Also MobileCLIP-B shows great scaling with number of seen samples (65.3$\to$71.7%) in comparison to previous work SLIP [42](42.8$\to$45.0%). In comparison to CLIPA [33] which uses multi-resolution training for efficiency, training with DataCompDR-12M is more efficient: CLIPA obtains 63.2% with 2.69B multi-resolution seen samples (which has equivalent compute as $\sim$0.5B $224^{2}$ seen samples), that is worse than MobileCLIP-B’s 65.3% with only 0.37B seen samples. Further, TinyCLIP-39M/16 in comparison to MobileCLIP-S2 has higher latency and less accuracy, and TinyCLIP-8M/16 is significantly less accurate than MobileCLIP-S0 (41.1% vs 59.1%) while having a close latency (2.6 ms vs 3.1 ms).
Name Dataset Seen Samples Latency (ms) (img+txt) Zero-shot IN-val
CLIP-B/16 [46, 42] CC-12M [4]
0.39B
11.5 + 3.3
36.5
CLIP-B/16 [46, 42] YFCC-15M [56]
0.37B
37.6
SLIP-B/16 [42] CC-12M [4]
0.39B
40.7
SLIP-B/16 [42] YFCC-15M [56]
0.37B
42.8
CLIP-B/16 DataComp-12M [18]
0.37B
10.4 + 3.3
50.1
MobileCLIP-B DataCompDR-12M
0.37B
10.4 + 3.3
65.3
CLIP-B/32 [46, 7] YFCC-15M [56]
0.49B
5.9 + 3.3
32.8
SLIP-B/32 [42, 7]
34.3
FILIP-B/32 [71, 7]
39.5
DeCLIP-B/32 [34]
43.2
DeFILIP-B/32 [7]
45.0
RILS-B/16 [70] LAION-20M [50]
0.5B
11.5 + 3.3
45.0
TinyCLIP-8M/16 [67] YFCC-15M [56]
0.75B
2.0 + 0.6
41.1
SLIP-B/16 [42] YFCC-15M [56]
0.75B
11.5 + 3.3
44.1
CLIP-B/16 DataComp-12M [18]
0.74B
10.4 + 3.3
53.5
MobileCLIP-S0 DataCompDR-12M
0.74B
1.5 + 1.6 59.1
TinyCLIP-39M/16 [67] YFCC-15M [56]
0.75B
5.2 + 1.9
63.5
MobileCLIP-S2 DataCompDR-12M
0.74B
3.6 + 3.3 64.6
MobileCLIP-B DataCompDR-12M
0.74B
10.4 + 3.3
69.1
SLIP-B/16 [42] YFCC-15M [56]
1.5B
11.5 + 3.3
45.0
CLIP-B/16 DataComp-12M [18]
1.48B
10.4 + 3.3
55.7
MobileCLIP-B DataCompDR-12M
1.48B
10.4 + 3.3
71.7
CLIPA-B/16 [33] LAION-400M [50]
${}^{\dagger}$
11.5 + 3.3
63.2
Table 5: Small-scale CLIP training. ${}^{\dagger}$ refers to multi-resolutions. Models are grouped based on the number of samples seen.
5.3 Learning Efficiency
Training longer with knowledge distillation is known to consistently improve performance for classification models [2]. In Fig. 5(a) we show our reinforced training also benefits from longer training, achieving 71.7% ImageNet-val zero-shot accuracy after 120 epochs using only a 12M subset of DataComp-1B. In comparison, non-reinforced training at best reaches 55.7% accuracy.
We also demonstrate scaling with dataset size in Fig. 5(b), where we deploy subsets of DataComp-1B from 1.28M to all 1.28B samples. For all experiments we train for 20k iterations with global batch size of 65k (equivalent to one epoch training on 1.28B subset). Training on DataCompDR reaches above 55.2% accuracy with 1.28M samples while training on DataComp-1B gets only to $\sim$6accuracy. In this setup, we observe more than 100$\times$ data efficiency using DataCompDR. Moreover, we observe 1000$\times$ data efficiency for performance on Flickr30k.
(a) Iteration Scaling
(b) Data Scaling
Figure 6: Learning efficiency up to 1000$\times$. Training on DataCompDR is 10$\times$ more iteration efficient and 100$\times$ more data efficient on ImageNet-val and 18$\times$ and 1000$\times$ more efficient on Flickr30k compared with non-reinforced training.
Name Dataset Seen Samples Image Encoder Text Encoder Params (M) (img+txt) Latency (ms) (img+txt) Zero-shot CLS Flickr30k Ret. COCO Ret. Avg. Perf. on 38
[…]
Ensemble Teacher
DataComp-1B [18]
OpenAI-400M [46]
ViT-L/14
ViT-L/14
Base
Base
80.1 69.6 74.5 92.3 46.7 66.5
67.3
TinyCLIP-RN19M [67] LAION-400M [50]
15.2B
ResNet-19M
Custom
18.6 + 44.8
1.9 + 1.9
56.3 43.6 58.0 75.4 30.9 47.8
48.3
TinyCLIP-RN30M [67] LAION-400M [50]
15.2B
ResNet-30M
Custom
29.6 + 54.2
2.6 + 2.6
59.1 45.7 61.5 80.1 33.8 51.6
50.2
TinyCLIP-40M/32 [67] LAION-400M [50]
15.2B
ViT-40M/32
Custom
39.7 + 44.5
3.0 + 1.9
59.8 46.5 59.1 76.1 33.5 48.7
51.2
MobileCLIP-S0 DataCompDR-1B
13B
MCi0
MCt
11.4 + 42.4
1.5 + 1.6
67.8 55.1 67.7 85.9 40.4 58.7 58.1
OpenAI-RN50 OpenAI-400M [46]
13B
ResNet-50
Base
38.3 + 63.4
3.3 + 3.3
59.8 45.1 57.4 80.0 28.5 48.8
48.1
TinyCLIP-61M/32 [67] LAION-400M [50]
15.2B
ViT-61M/32
Custom
61.4 + 54.0
4.3 + 2.6
62.4 48.7 62.6 78.7 36.5 52.8
53.0
TinyCLIP-63M/32 [67]
LAION-400M [50]
YFCC-15M [56]
15.8B
ViT-63M/32
Custom
64.5 (-) 66.0 84.9 38.5 56.9
MobileCLIP-S1 DataCompDR-1B
13B
MCi1
Base
21.5 + 63.4
2.5 + 3.3
72.6 60.7 71.0 89.2 44.0 62.2 61.3
OpenAI-RN101 OpenAI-400M [46]
13B
ResNet-101
Base
56.3 + 63.4
4.3 + 3.3
62.3 48.5 58.0 79.0 30.7 49.8
50.3
OpenAI-B/32 OpenAI-400M [46]
13B
ViT-B/32
Base
86.2 + 63.4
5.9 + 3.3
63.3 48.5 58.8 78.9 30.4 50.1
52.5
LAION-B/32 LAION-2B [51]
32B
65.7 51.9 66.4 84.4 39.1 56.2
54.8
DataComp-B/32 DataComp-1B [18]
13B
69.2 55.2 61.1 79.0 37.1 53.5
58.0
DataComp-B/32-256 DataComp-1B [18]
34B
ViT-B/32-256
Base
86.2 + 63.4
6.2 + 3.3
72.8 58.7 64.9 84.8 39.9 57.9
60.9
MobileCLIP-S2 DataCompDR-1B
13B
MCi2
Base
35.7 + 63.4
3.6 + 3.3
74.4 63.1 73.4 90.3 45.4 63.4 63.7
VeCLIP-B/16 [31] WIT-200M
6.4B
ViT-B/16
Base
86.2 + 63.4
11.5 + 3.3
64.6 (-) 76.3 91.1 48.4 67.2
OpenAI-B/16 WIT-400M [46]
13B
Base
86.2 + 63.4
11.5 + 3.3
68.3 55.9 67.7 85.9 40.4 58.7
58.1
LAION-B/16 LAION-2B [51]
34B
Base
86.2 + 63.4
11.5 + 3.3
70.2 56.6 69.8 86.3 42.3 59.4
58.7
EVA02-B/16 Merged-2B [54]
8B
Base
86.2 + 63.4
74.7 59.6 71.5 86.0 42.2 58.7
58.9
DFN-B/16 DFN-2B [16]
13B
Base
86.2 + 63.4
11.5 + 3.3
76.2 62.3 69.1 85.4 43.4 60.4
60.9
DataComp-B/16 DataComp-1B [18]
13B
Base
86.2 + 63.4
11.5 + 3.3
73.5 60.8 69.8 86.3 42.3 59.4
61.5
SigLIP-B/16 [76] Webli-1B
40B
Custom
92.9 + 110.3
9.9 + 5.8
76.0 61.0 74.7 89.1 47.8 65.7
62.3
MobileCLIP-B DataCompDR-1B
13B
Base
86.3 + 63.4
10.4 + 3.3
76.8 65.6 77.3 91.4 50.6 68.8 65.2
Table 6: MobileCLIP family of models has the best average performance at various latencies. Retrieval performances are reported @1. Last column shows average performance on 38 datasets as in OpenCLIP [28]. Models are grouped by their total latency in increasing order and by performance within each group. “Base” refers to standard CLIP Transformer-based [62] text encoder with 12 layers, and “Custom” stands for customized text encoder used in the respective method. For TinyCLIP-63M/32 and EVA02-B/16, we were unable to reliably benchmark models. Note: EVA02-B/16 [54] uses MIM pretrained weights for its vision encoder and OpenCLIP-B pretrained weights for its text encoder. TinyCLIP models use advanced weight initialization methods utilizing OpenCLIP models trained on LAION-2B[51] dataset. All other models, including ours are trained from scratch.
5.4 Comparison with State-of-the-art
In Tab. 6, we compare with methods with large scale training. MobileCLIP-S0, trained on DataCompDR-1B significantly outperforms recent works like TinyCLIP [67], and has similar performance as a ViT-B/32 model trained on DataComp [18] while being 2.8$\times$ smaller and 3$\times$ faster. MobileCLIP-S2 obtains 2.8% better average performance on 38 datasets and significantly better retrieval performance when compared to ViT-B/32-256 model trained 2.6$\times$ longer on DataComp [18]. MobileCLIP-S2 is 1.5$\times$ smaller and 1.4$\times$ faster than ViT-B/32-256 model. MobileCLIP-B obtains 2.9% better average performance on 38 datasets and better retrieval performance while being 26.3% smaller than SigLIP-B/16 [76] model, which is trained approximately 3$\times$ longer on WebLI dataset.
5.5 Retrieval Performance Analysis
We evaluate our models on the recent Attribute, Relation and Order (ARO) benchmark [74]. We compare our MobileCLIP-B trained on DataCompDR-1B with all the publicly available ViT-B/16:Base models in Tab. 7. Optimizing solely for zero-shot classification or retrieval using noisy webscale datasets can degrade the compositional understanding of natural scenes. DataCompDR largely improves the models performance on ARO benchmark while obtaining good performance on zero-shot classification and retrieval tasks. Compared to the recent SigLIP method [76], MobileCLIP-B obtains 19.5% and 12.4% better accuracy on Visual Genome Relation and Attributes datasets and achieves improved recall@1 on Flickr30k-Order and COCO-Order datasets by 69.7% and 50.3%, respectively.
Method Dataset IN-val VG VG COCO Flickr30k
zero-shot Rel. Attr. Order Order
CLIP OpenAI-400M [46] 68.3 58.7 62.2 50.4 57.3
CLIP LAION-2B [51] 70.2 39.7 62.3 31.0 37.5
CLIP DataComp-1B [18] 73.5 35.9 57.0 29.6 35.2
SigLIP [76] Webli-1B 76.0 35.1 56.0 32.7 40.7
CLIP DFN-2B [16] 76.2 33.1 57.4 18.5 22.5
MobileCLIP-B DataCompDR-1B 76.8 54.6 68.4 55.5 61.2
Table 7: Performance on ARO benchmark. All the models use ViT-B/16 as image encoder and the standard CLIP text encoder. For VG Rel. and VG Attr. datasets, Macro Accuracy is reported and for Flickr30k-Order and COCO-Order recall@1 is reported. For more details please refer to [74].
6 Conclusion
In this work we introduced MobileCLIP aligned image-text backbones, designed for on-device CLIP inference (low latency and size). We also introduced DataCompDR, a reinforcement of DataComp with knowledge from a pre-trained image captioning model and an ensemble of strong CLIP models. We demonstrated 10$\times$-1000$\times$ learning efficiency with our reinforced dataset. MobileCLIP models trained on DataCompDR obtain state-of-the-art latency-accuracy tradeoff when compared to previous works. MobileCLIP models also exhibit better robustness and improved performance on Attribute, Relation and Order (ARO) benchmark.
Supplementary Material
Appendix A Experimental Setup
Additional details of our training and evaluation are provided here. We train all models at 224 resolution.
Table 10 summarizes the hyperparameters we used to train MobileCLIP-B on DataCompDR-1B. For other variants of MobileCLIP (S0, S1, and S2) we use the same hyperparameters except using $\lambda=1.0$. For experiments on DataCompDR-12M we use global batch size of 8192.
For our ensemble distillation ablations in Appendix C, we use 32 total A100 GPUs but we use the same global batch size of 8192 as our other ablations. We also use a smaller uniformly sampled DataComp-8M for ablations in Appendices C and B that results in a slightly lower performance than DataCompDR-12M used for the rest of ablations.
For ImageNet-1k experiments in Sec. 4.2, we follow the training recipe prescribed in [58, 37], i.e. the models are trained for 300 epochs using AdamW optimizer with weight decay of 0.05 and peak learning rate $10^{-3}$ for a total batch size of 1024. The number of warmup epochs is set to 5 and cosine schedule is used to decay the learning rate. The teacher model for distillation is RegNetY-16GF [47] Our implementation uses Timm library [65] and all the models were trained on single machine with 8$\times$×NVIDIA A100 GPUs. The hyperparameters for the three variants of MCi are detailed in Tab. 8. The performance of MCi variants is detailed in Tab. 9 and compared against recent state-of-art efficient architectures. MCi obtains the best trade-off amongst recent efficient architectures as seen in Fig. 7.
Hyperparameter Training
MCi0, MCi1, MCi2
Stochastic depth rate [0.0, 0.05, 0.15]
Input resolution 256$\times$256
Data augmentation RandAugment
Mixup $\alpha$ 0.8
CutMix $\alpha$ 1.0
Random erase prob. 0.25
Label smoothing 0.1
Train epochs 300
Warmup epochs 5
Batch size 1024
Optimizer AdamW
Peak learning rate 1e-3
LR. decay schedule cosine
Weight decay rate 0.05
Gradient clipping
EMA decay rate 0.9995
Table 8: Training hyperparameters for ImageNet-1k experiments.
Model Eval Param FLOPs Mobile Top-1
Image Latency Acc.
Size (M) (G) (ms) (%)
MobileViG-M [43] 224 14.0 1.5 1.4 80.6
SwiftFormer-L1 [52] 224 12.1 1.6 1.5 80.9
EfficientFormerV2-S2 [37] 224 12.6 1.3 1.6 81.6
FastViT-SA12 [61] 256 11.5 1.9 1.5 81.9
MCi0 (ours) 256 11.8 2.4 1.5 82.2
MobileViG-B [43] 224 26.7 2.8 2.3 82.6
SwiftFormer-L3 [52] 224 28.5 4.0 2.6 83.0
EfficientFormerV2-L [37] 224 26.1 2.6 2.6 83.3
FastViT-SA24 [61] 256 21.5 3.8 2.4 83.4
MCi1 (ours) 256 21.9 4.7 2.5 83.8
FastViT-MA36 [61] 256 43.9 7.8 4.3 84.5
MCi2 (ours) 256 36.3 7.8 3.6 84.5
Table 9: Comparison of MCi variants with recent state-of-the-art models on ImageNet classification task.
Figure 7: Top-1 Accuracy on ImageNet v/s latency plot of MCi variants and recent state-of-the-art architectures.
Hyperparameter Value
MobileCLIP-B, S0, S1, S2
Input resolution ${}^{2}$, ${}^{2}$, ${}^{2}$, ${}^{2}$
Context length 77
Data augmentation RandAugment
Random resize crop scale [0.08, 1.0]
Random resized crop ratio [0.75, 1.33]
RangeAugment target value (40, 20)
Train iterations 200k
Warmup iterations 2k
Global batch size 65536
Optimizer AdamW
AdamW beta1 0.9
AdamW beta2 0.95
Max learning rate 1e-3
Min learning rate 1e-6
LR. decay schedule cosine
Weight decay rate 0.2
Gradient clipping
Mixed precision BFloat16
EMA decay rate 0.9995
CLIP loss weight 0.25
KD loss weight 0.75
GT caption weight 1.0
Synth. caption weight 1.0
Synth. teacher coca_ViT-L-14
Teacher 1 openai-ViT-L-14
Teacher 2 datacomp_xl_s13b_b90k-ViT-L-14
Teacher resolution 224$\times$224
Table 10: Training hyperparameters for our CLIP experiments on DataCompDR.
Appendix B Image Augmentation
In this section we provide a detailed ablation on the effect of image augmentations. The training setup is the same as training with DataCompDR-12M presented in Sec. 5.2, except we used an 8M subset for this ablation. In Tab. 11 we show classification and retrieval performance of a ViT-B/16 based CLIP model trained with our final loss as in Eq. 3 ($\lambda=1$) and different image augmentations. Note that we feed the same augmented image to both teacher and student models. First, we consider RandomResizedCrop (RRC) with three magnitudes (0.08, 0.4, 0.9) determining the lower bound of random area of the crop (smaller lower bound means stronger augmentation). We observe that strong RRC results in significant accuracy improvement both for classification and retrieval metrics. While using strong RRC augmentation is standard for supervised training, for CLIP training the widely used recipe [46] includes weak RRC (lower-bound for scale= 0.9).
We further utilize RangeAugment [41] to automatically adjust Brightness, Contrast, and Noise. We use PSNR metric with target range [20, 40] and a Cosine curriculum. Since in RangeAugment individual augmentation magnitudes are adjusted dynamically during training, they cannot be stored as part of the dataset reinforcement process. Hence, we only apply it to images fed to the student model. We show that if the same augmentation is applied to both student and teacher (not feasible for our dataset reinforcement approach) further improvement can be obtained (56.6% vs 55.9% on ImageNet-val).
Finally, we consider RandomHorizontalFlip, RandomErasing [77], and RandAugment [6], and find that only RandAugment is beneficial in our setup. Our reinforced datasets include parameters of RRC and RandAugment and during training time we apply RangeAugment to images fed to the student model.
Image Augmentations Zero-shot CLS Flickr30k Ret. COCO Ret. Avg Perf.
IN-val IN-shift I2T T2I I2T T2I on 38
RandomResizedCrop: 0.9-1.0
Student-RangeAugment [41]
51.0 40.1 54.2 68.5 30.5 45.3 45.9
RandomResizedCrop: 0.4-1.0
Student-RangeAugment
55.0 43.9 60.4 76.0 34.1 48.4 48.9
RandomResizedCrop: 0.08-1.0
Student-RangeAugment
55.9 44.6 58.8 76.1 34.2 49.0 49.6
RandomResizedCrop: 0.08-1.0
56.4 44.6 59.8 74.6 34.4 49.3 49.1
RandomResizedCrop: 0.08-1.0
Student&Teacher-RangeAugment
56.6 44.9 60.2 74.0 34.9 50.5 50.8
RandomResizedCrop: 0.08-1.0
Student-RangeAugment
RandomHorizontalFlip: p=0.5
55.9 44.7 59.4 75.9 34.4 49.2 48.8
RandomResizedCrop: 0.08-1.0
Student-RangeAugment
RandomErasing [77]: p=0.25
55.8 44.5 59.4 75.3 34.5 49.7 49.1
RandomResizedCrop: 0.08-1.0
Student-RangeAugment
RandAugment [6]
56.6 45.4 60.9 78.3 35.0 51.0 50.2
Table 11: Ablation on different augmentations for distillation. We highlight our choice with blue.
Appendix C CLIP Ensembles
In this section we provide a detailed ablation on CLIP ensembles. First, we show that we can construct more accurate zero-shot models by ensembling pretrained individual CLIP models. For inference, we concatenate normalized embeddings of each modality followed by a re-normalization. In Tab. 12 we show performance of some CLIP ensemble models that we picked from OpenCLIP [28]. We also include performance of individual models. Evidently, ensembling results in improved performance. For example, an ensemble of two pretrained ViT-L-14-based CLIP models from datacomp_xl_s13b_b90k and openai results in average performance of 67.3%, while each individual model has 66.3% and 61.7% performance, respectively. Further, ensembling can be a more parameter efficient approach to obtain a stronger model. For instance, the ensemble of two ViT-L-14-based CLIP models has less parameters than the one with ViT-bigG-14 image encoder, but comes with the same ImageNet-val performance (80.1%). In general, given a set of pretrained CLIP models (e.g., as in OpenCLIP [28]) with this approach we can push state-of-the-art and obtain stronger zero-shot performance. Here, we show and ensemble of four CLIP models can reach up to 81.7% zero-shot classification performance on ImageNet-val, while individual models’ performance is not more than 80.1%. As stronger individual models become publicly available, one can create stronger ensembles with this approach.
In this work, we are interested in creating a strong ensemble model to be used as a teacher in the context of distillation. In Tab. 13 we show performance of a ViT-B/16 CLIP model trained with different CLIP models as teacher (both individual models and ensembles). Training setup is the same as that of in Sec. 5.2, except we use a uniformly sampled 8M subset. Similar to standard distillation for classification task [25], we observe that more accurate CLIP models are not necessarily better teachers. We picked the ensemble of two ViT-L-14-based CLIP models as the teacher model (highlighted in blue) in our dataset reinforcement process.
Teacher Teacher Teacher Zero-shot CLS Flickr30k Ret. COCO Ret. Avg Perf.
Models(s) Pre-taining(s) Resolution(s) IN-val IN-shift I2T T2I I2T T2I on 38
ViT-bigG-14 laion2b_s39b_b160k 224 80.1 69.1 79.6 92.9 51.4 67.4 66.7
EVA01-g-14-plus merged2b_s11b_b114k 224 79.3 69.3 79.0 91.7 50.3 68.2 66.2
ViT-L-14 datacomp_xl_s13b_b90k 224 79.2 67.9 73.4 89.0 45.7 63.3 66.3
ViT-L-14 openai 224 75.5 64.9 65.0 85.2 36.5 56.3 61.7
ViT-L-14-336 openai 336 76.6 67.1 66.9 87.7 37.1 57.9 62.8
ViT-L-14
ViT-L-14
datacomp_xl_s13b_b90k
openai
224
224
80.1 69.6 74.5 92.3 46.7 66.5 67.3
ViT-L-14
ViT-L-14-336
datacomp_xl_s13b_b90k
openai
224
336
80.5 70.6 75.8 91.8 47.0 67.0 67.8
EVA01-g-14-plus
ViT-L-14
ViT-L-14
merged2b_s11b_b114k
datacomp_xl_s13b_b90k
openai
224
224
224
81.1 70.9 78.1 93.8 50.2 69.7 68.5
EVA01-g-14-plus
ViT-L-14
ViT-L-14-336
merged2b_s11b_b114k
datacomp_xl_s13b_b90k
openai
224
224
336
81.2 71.6 78.8 93.7 50.2 69.9 68.9
convnext_xxlarge
ViT-L-14
ViT-L-14-336
laion2b_s34b_b82k_augreg_soup
datacomp_xl_s13b_b90k
openai
256
224
336
81.5 71.7 79.0 94.5 50.5 69.5 68.7
ViT-bigG-14
EVA01-g-14-plus
ViT-L-14
ViT-L-14
laion2b_s39b_b160k
merged2b_s11b_b114k
datacomp_xl_s13b_b90k
openai
224
224
224
224
81.6 71.7 79.9 94.6 52.4 71.3 69.4
EVA01-g-14-plus
ViT-L-14-336
ViT-L-14
convnext_xxlarge
merged2b_s11b_b114k
openai
datacomp_xl_s13b_b90k
laion2b_s34b_b82k_augreg_soup
224
336
224
256
81.7 72.1 80.0 95.0 52.0 70.8 69.3
ViT-L-14
ViT-L-14-336
RN50x64
RN50x16
openai
openai
openai
openai
224
336
384
448
78.2 68.9 73.4 89.7 42.0 63.5 65.5
Table 12: Zero-shot evaluation of (ensemble of) clip models. Each group of rows corresponds to an ensemble teacher. All models are taken from OpenCLIP [28] on Aug-2023. We highlight our choice with blue.
Teacher Teacher Teacher Zero-shot CLS Flickr30k Ret. COCO Ret. Avg Perf.
Models(s) Pre-taining(s) Resolution(s) IN-val IN-shift I2T T2I I2T T2I on 38
ViT-bigG-14 laion2b_s39b_b160k 224 53.4 42.6 59.6 76.2 35.8 52.1 47.8
EVA01-g-14-plus merged2b_s11b_b114k 224 54.5 43.3 59.6 74.6 35.4 50.8 47.7
ViT-L-14 datacomp_xl_s13b_b90k 224 54.0 43.4 58.9 74.3 34.3 50.1 48.3
ViT-L-14 openai 224 54.4 42.7 54.5 69.1 29.7 44.6 47.2
ViT-L-14-336 openai 336 54.2 43.3 53.6 68.7 30.1 44.3 47.2
ViT-L-14
ViT-L-14
datacomp_xl_s13b_b90k
openai
224
224
56.3 44.8 59.2 74.5 34.4 49.9 49.6
ViT-L-14
ViT-L-14-336
datacomp_xl_s13b_b90k
openai
224
336
55.9 44.6 58.8 76.1 34.2 49.0 49.6
EVA01-g-14-plus
ViT-L-14
ViT-L-14
merged2b_s11b_b114k
datacomp_xl_s13b_b90k
openai
224
224
224
56.2 45.0 59.6 76.9 35.7 51.5 49.4
EVA01-g-14-plus
ViT-L-14
ViT-L-14-336
merged2b_s11b_b114k
datacomp_xl_s13b_b90k
openai
224
224
336
56.0 44.5 60.1 76.5 35.3 50.6 49.5
convnext_xxlarge
ViT-L-14
ViT-L-14-336
laion2b_s34b_b82k_augreg_soup
datacomp_xl_s13b_b90k
openai
256
224
336
55.8 44.4 59.4 75.1 35.0 49.5 50.1
ViT-bigG-14
EVA01-g-14-plus
ViT-L-14
ViT-L-14
laion2b_s39b_b160k
merged2b_s11b_b114k
datacomp_xl_s13b_b90k
openai
224
224
224
224
56.3 44.6 60.8 76.2 35.8 51.4 49.2
EVA01-g-14-plus
ViT-L-14-336
ViT-L-14
convnext_xxlarge
merged2b_s11b_b114k
openai
datacomp_xl_s13b_b90k
laion2b_s34b_b82k_augreg_soup
224
336
224
256
55.9 44.6 60.4 75.1 35.6 52.3 49.4
ViT-L-14
ViT-L-14-336
RN50x64
RN50x16
openai
openai
openai
openai
224
336
384
448
56.4 44.6 57.9 72.0 31.7 47.0 48.6
Table 13: Ablation on using different (ensemble of) teacher models in our multi-modal distillation. Each group of rows demonstrate an ensemble teacher. Student architecture is fixed to ViT-B/16 for image encoder and base 12-layer Transformer for text encoder (MobileCLIP-B setup). For this ablation, we use an 8M subset of DataComp and train all experiments for 20k iterations with global batch size of 8k. All models are imported from OpenCLIP [28] on Aug-2023. We highlight our choice with blue.
Appendix D Ablations on Lossy Compressions
In Tab. 2(c) we presented the storage sizes for DataCompDR-12M and DataCompDR-1B with BFloat16 compression of the embeddings. In this section, we further analyze the storage reduction by i) reducing the number of augmentations, and ii) lossy compression of embeddings.
We report the total storage size for 12.8k samples of DataCompDR in Tab. 14. The storage size for DataCompDR-12M can be easily deduced by multiplying the numbers by 1000 (TBs instead of GBs) and by $10^{5}$ for DataCompDR-1B.
Table 15 shows the accuracy of training with BFloat16 embeddings achieves accuracies within the standard deviation of the training on DataComp-12M.
Image Text Syn.
Aug. Params
Text Emb.
Image Emb.
BFloat16
Sparsity
Size (GBs)
0.9
0.9
5+1
30
3.3
5+1
30
1.9
5+1
30
50%
1.8
5+1
30
50%
1.3
5+1
10
1.9
5+1
10
1.4
5
5
1.5
5
5
1.2
2
2
1.1
2
2
1.0
Table 14: Total storage for 12.8k samples stored in individual Pickle Gzip files. Storage for 12.8M and 1.28B samples are approximately the same numbers in TBs and 100 TBs.
Num. Aug. 1 2 5 10 15 20 25 30
w/o BFloat16 60.63 63.27 64.81 64.74 64.49 64.92 64.78 64.74
w/ BFloat16 - - 64.32 64.88 64.57 64.81 65.13 64.91
Table 15: Effect of BFloat16 and the number of augmentations on ImageNet-val zero-shot Accuracy. We train on DataCompDR-12M for approximately 30 epochs.
Appendix E Hybrid Text Encoder
In this section, we ablate over kernel dimensions for our hybrid text encoder. For this ablation, we use a 6-layered fully convolutional text encoder and systematically increase the kernel size. We use ViT-B/16 as the image encoder for these runs. These models were trained on DataCompDR-12M for 30k iterations. From Tab. 16, we notice that zero-shot IN-val performance does improve with increased kernel size, but it is significantly more expensive to run the model on mobile device. For zero-shot IN-val performance improvement of 1.1%, the model is 4.5$\times$ slower. From Tab. 16, kernel size of 11 obtains the best accuracy-latency trade-off.
Kernel Size 3 11 31
Num Params. (M) 38.2 38.3 38.4
Latency (ms) 1.0 1.2 5.4
IN-val 56.3 57.9 59.0
Table 16: Ablation on kernel size for text encoder. We train for 30k iterations. We highlight our choice with blue
Appendix F Extended Results
In this section we provide extended zero-shot results of our proposed family of CLIP models: MobileCLIP-S0, MobileCLIP-S1, MobileCLIP-S2, and MobileCLIP-B. Zero-shot classification and retrieval results are provided in Tab. 17. We also include additional results from related works where only partial evaluation is available.
Name ImageNet Shifts CLS Flickr30k Retrieval COCO Retrieval
[…]
@1 @5 @10 @1 @5 @10 @1 @5 @10 @1 @5 @10
MobileCLIP-B 76.8 58.7 89.6 41.4 64.5 69.8 69.4 91.4 99.1 99.9 77.3 94.4 96.7 68.8 88.3 92.9 50.6 74.9 82.9
MobileCLIP-S2 74.4 49.3 87.0 46.9 62.2 66.8 66.6 90.3 98.9 99.6 73.4 92.3 95.6 63.4 85.1 91.4 45.4 70.1 79.0
MobileCLIP-S1 72.6 40.3 84.7 50.5 60.3 64.9 63.4 89.2 98.0 99.5 71.0 91.3 95.3 62.2 84.3 90.1 44.0 68.9 77.7
MobileCLIP-S0 67.8 26.5 78.6 53.8 55.5 59.9 55.9 85.9 97.1 98.8 67.7 88.8 93.3 58.7 81.1 88.2 40.4 66.0 75.9
DIME-FM-B/32 [55] 66.5 32.2 69.8 (-) 46.5 58.9 43.2 (-) (-) (-) (-) (-) (-) (-) (-) (-) (-) (-) (-)
VeCLIP-B/16 [31] 64.6 (-) (-) (-) (-) 57.7 (-) 91.1 98.5 99.7 76.3 93.5 96.4 67.2 87.3 92.7 48.4 73.3 81.8
TinyCLIP-63M/32 [67] 64.5 22.8 74.1 (-) 50.8 55.7 31.2 84.9 (-) (-) 66.0 (-) (-) 56.9 (-) (-) 38.5 (-) (-)
CLIPA-B/16 [33] 63.2 26.8 73.2 (-) 48.7 55.6 44.3 75.3 (-) (-) 58.3 (-) (-) 53.1 (-) (-) 35.2 (-) (-)
Table 17: Extended zero-shot evaluations. We also include additional results from related works where the full DataComp [18] evaluation was not accessible. Numbers are read from the corresponding papers. For each method we picked their best model up to ViT-B/16 size. Please see Tab. 6 for additional details including runtime benchmarking. Models are sorted by their zero-shot classification performance on ImageNet-val. Here our MobileCLIP-S1 is fully trained with 13B seen samples.